OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

What are your unknowns in this problem? Can you write any mathematical relationships based on your unknowns and what you know?

OpenStudy (anonymous):

wait i think i got it

OpenStudy (anonymous):

Okay. What did you get? I'll see if I got the same!

OpenStudy (anonymous):

If HM represents HUMAN and HR represents HORSE then...

OpenStudy (anonymous):

HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50 24 horses and 50 humans

OpenStudy (anonymous):

Looks good to me! Nice work.

OpenStudy (anonymous):

That's a textbook algebra solution. :)

OpenStudy (anonymous):

Thanks